# Robust Point Matching with Distance Profiles

### v5

@@ 1.2 Stylized Example
Figure 1(c) and Figure 1(f) repeat the same procedure after removing 16 points around the fin of the target dolphin in Figure 1(e) and clustering again with $k=6$. Figure 1(c) shows that the matching result is similar to the previous one. In this case, though the points around the fin of the source dolphin do not have corresponding points in the target dolphin and thus we may ignore matching them, the proposed procedure still matches them to the points around the peduncle (magenta) of the target dolphin, which is a cluster adjacent to the removed fin. This stylized example confirms the performance of the proposed distance profile matching procedure in practice.
2 Outlier Robustness of Distance Profile Matching
This section analyzes the outlier robustness of the distance profile matching procedure described in Algorithm 1. As we have seen in the previous section, Algorithm 1 is designed to match the points corresponding to the same region of the objects, while ignoring or arbitrarily matching the points from the outlier parts. In order to theoretically analyze such outlier robustness, we consider appropriate model assumptions on the underlying objects $\mu,\nu\in\mathscr{P}(\mathbb{R}^{d})$ that capture the nature of the outliers. Motivated by several practical considerations, we model $\mu,\nu$ as follows.
Assumption 1.
$\mu=\sum_{k=1}^{t}p_{k}N(\theta_{k},\sigma_{k}^{2}I_{d})\quad\text{and}\quad\nu=\sum_{k=1}^{s}q_{k}N(\eta_{k},\tau_{k}^{2}I_{d}),$
For some $K\leq\min\{t,s\}$, we assume that the first $K$ components of $\mu$ and $\nu$ are identical, that is, $p_{k}=q_{k}$, $\theta_{k}=\eta_{k}$, and $\sigma_{k}=\tau_{k}$ for $k\in[K]$.
Assumption 1 is equivalent to letting $\mu=\lambda\rho+(1-\lambda)\mu_{o}$ and $\nu=\lambda\rho+(1-\lambda)\nu_{o}$, where the common part $\lambda\rho$ is the mixture consisting of the first $K$ components and the other parts, namely, the outlier parts $(1-\lambda)\mu_{o},(1-\lambda)\nu_{o}$ are the mixtures of the remaining components. The motivation behind Assumption 1 is that the objects $\mu,\nu$ are composed of several regions as we have seen from the example in Section 1.1, where each region is represented as a Gaussian distribution so that the mean corresponds to the center of the region and the variance roughly represents the size of the region. The points corresponding to the same region are to be matched, while we ignore matching the points from the outlier components.
$\bigcap_{i\in[n]:t(i)\in[K]}(s(\pi(i))=t(i)),$ (2.1)
Theorem 1.
Suppose $\mu$ and $\nu$ satisfy Assumption 1. Let $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ be independent samples from $\mu$ and $\nu$, respectively. Let $\pi\colon[n]\to[m]$ be the output of Algorithm 1 with $\mathcal{D}=W_{1}$. For $i\in[n]$ and $j\in[m]$, let $t(i)\in[t]$ and $s(j)\in[s]$ denote the components where $X_{i}$ and $Y_{j}$ are from, respectively. Then, for any $\delta\in(0,1)$, we have
if
$\begin{split}&\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)-R\cdot\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)}\\ &\geq\left(16\Gamma\sqrt{d}\right)\vee\bigg{(}8R\sqrt{\frac{\log K+\log(t+s-2)+\log\tfrac{4}{\delta}}{2(n\wedge m)}}\bigg{)}\vee\bigg{(}32\Gamma\sqrt{2\log\frac{4(n\vee m)^{2}}{\delta}}\bigg{)},\end{split}$ (2.3)
where $R=\max\{\|x-y\|_{2}:x,y\in\{\theta_{1},\ldots,\theta_{t},\eta_{1},\ldots,\eta_{s}\}\}$, $\Gamma=\max_{k\in[t]}\sigma_{k}\vee\max_{k\in[s]}\tau_{k}$, and
To understand Theorem 1, it is instructive to begin with the clean case where the noise is absent, that is, $\Gamma=0$, and there is no sampling error, namely, exactly $np_{k}$ and $mq_{k}$ points among $X_{i}$’s and $Y_{j}$’s are from the $k$-th component of $\mu$ and $\nu$, respectively. Then, the point clouds are essentially the sets of centers $\{\theta_{1},\ldots,\theta_{t}\}$ and $\{\eta_{1},\ldots,\eta_{s}\}$ weighted by $(p_{1},\ldots,p_{t})$ and $(q_{1},\ldots,q_{s})$, respectively. Therefore, the Wasserstein-1 distances between distance profiles (1.1) boil down to the Wasserstein-1 distance between the weighted sets of centers (2.4). Then, the success of Algorithm 1, namely, the event (2.1), is equivalent to having $\operatorname*{arg\,min}_{\beta\in[s]}\overline{W}(\alpha,\beta)=\alpha$ for any $\alpha\in[K]$, which is guaranteed by having
$\min_{\beta\in[s]\backslash\{\alpha\}}\overline{W}(\alpha,\beta)>\overline{W}(\alpha,\alpha).$ (2.5)
To ensure this for all $\alpha\in[K]$ simultaneously, we require a stronger condition, [^2] Note that (2.6) is much stronger than requiring (2.5) for all $\alpha\in[K]$. We chose to present Theorem 1 based on (2.6) as it leads to a more interpretable condition in terms of the outlier proportion. namely,
$\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)-\max_{\alpha\in[K]}\overline{W}(\alpha,\alpha)>0.$ (2.6)
How can this condition be satisfied? First, suppose there is no outlier at all, namely, $K=t=s$. Then, since $\overline{W}(\alpha,\alpha)=0$ for any $\alpha\in[K]$, it suffices to have $\overline{W}(\alpha,\beta)>0$ for any $\beta\in[s]\backslash\{\alpha\}$, which can be guaranteed if the pairwise distances among centers are well-separated. In general, we can upper bound $\overline{W}(\alpha,\alpha)$ for any $\alpha\in[K]$, which is essentially controlled by the outlier proportion $1-\sum_{k\in[K]}p_{k}$, as we show in Lemma 1. Meanwhile, for $\alpha\in[K]$ and $\beta\in[s]\backslash\{\alpha\}$, the quantity $\overline{W}(\alpha,\beta)$ is essentially of a constant order, namely, $O(1)$. Therefore, the condition (2.6) is satisfied if the outlier proportion is small enough, which is captured by the left-hand side of (2.3).
In general cases with $\Gamma>0$, where the point clouds are samples, the Wasserstein-1 distances (1.1) deviate from (2.4) due to the noise and sampling error. Therefore, the condition (2.6) needs to be modified to reflect this deviation, and characterizing this probabilistic deviation leads to the right-hand side of (2.3).
Based on the above insights, we prove Theorem 1 as follows.
@@ 2 Outlier Robustness of Distance Profile Matching
Meanwhile, notice that $X_{i}=\theta_{t(i)}+\xi_{i}$ and $Y_{j}=\eta_{s(j)}+\zeta_{j}$ for any $(i,j)\in[n]\times[m]$, where $\xi_{i}\,|\,t(i)\sim N(0,\sigma_{t(i)}^{2}I_{d})$ and $\zeta_{j}\,|\,s(j)\sim N(0,\tau_{s(j)}^{2}I_{d})$ are independent. Recall that the Wasserstein-1 distance between two sets of points in $\mathbb{R}$ with the same cardinality satisfies the following:
@@
$\begin{split}&\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}W(i,j)-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}W(i,j)\\ &\geq\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}\overline{W}(t(i),s(j))-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}\overline{W}(t(i),s(j))-2R(Z_{1}+Z_{2})-2(U+V)\\ &\geq\min_{\beta\in[s]\backslash\{t(i)\}}\overline{W}(t(i),\beta)-\overline{W}(t(i),t(i))-2R(Z_{1}+Z_{2})-2(U+V)\\ &\geq\underbrace{\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)-R\cdot\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)}}_{=:\Omega}-2R(Z_{1}+Z_{2})-2(U+V),\end{split}$ (2.7)
@@
Meanwhile, for $i\neq\ell$, note that we have $\xi_{i}-\xi_{\ell}\,|\,t(i),t(\ell)\sim N(0,(\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2})I_{d})$, which implies $\|\xi_{i}-\xi_{\ell}\|_{2}^{2}\,|\,t(i),t(\ell)\sim(\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2})\chi^{2}(d)$. Hence, we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq\sum_{i\neq\ell}\mathbb{P}\left(\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{(\Omega/8)^{2}}{2\sigma^{2}}\right).$
As $\Omega^{2}\geq 256d\cdot\max_{k\in[t]}\sigma_{k}^{2}$, we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{(\Omega/8)^{2}}{2\sigma^{2}}\right)\leq n^{2}\exp\left(-\frac{(\Omega/8)^{2}}{32\sigma^{2}}\right).$
Similarly, as $\Omega^{2}\geq 256d\cdot\max_{k\in[s]}\tau_{k}^{2}$, we have
$\mathbb{P}\left(\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq m^{2}\exp\left(-\frac{(\Omega/8)^{2}}{32\tau^{2}}\right).$
@@
$\begin{split}\mathbb{P}(E)&\leq\mathbb{P}\left(Z_{1}\geq\frac{\Omega}{4R}\right)+\mathbb{P}\left(Z_{2}\geq\frac{\Omega}{4R}\right)+\mathbb{P}\left(U\geq\frac{\Omega}{8}\right)+\mathbb{P}\left(V\geq\frac{\Omega}{8}\right)\\ &\leq 2K(t+s-2)\exp\left(-\frac{2(n\wedge m)\Omega^{2}}{(8R)^{2}}\right)+2(n\vee m)^{2}\exp\left(-\frac{(\Omega/8)^{2}}{32\Gamma^{2}}\right)\\ &\leq\frac{\delta}{2}+\frac{\delta}{2}=\delta,\end{split}$
@@
$\Omega_{o}:=\min_{\alpha\in[K]}\left(\min_{\beta\in[s]\backslash\{\alpha\}}\overline{W}(\alpha,\beta)-\overline{W}(\alpha,\alpha)\right).$
To see this, it suffices to observe that we can replace $\Omega$ in the last line of (2.7) with $\Omega_{o}$, and the rest of the proof follows the same with $\Omega_{o}$ in place of $\Omega$. Since $\Omega_{o}\geq\Omega$, the condition (2.3) is a stronger condition than requiring $\Omega_{o}$ to be at least the right-hand side of (2.3). As we briefly mentioned before the proof, we chose to present (2.3) as it is written in terms of the outlier proportion which is more interpretable.
@@
$\overline{W}(\alpha,\alpha)\leq\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)}\cdot\max_{(\alpha,\beta)\in[t]\times[s]}\|\theta_{\alpha}-\eta_{\beta}\|_{2}.$ (2.8)
@@
where $\lambda=\sum_{k\in[K]}p_{k}$. Since $|\|\theta_{\alpha}-\theta_{k}\|_{2}-\|\theta_{\alpha}-\eta_{\ell}\|_{2}|\leq\|\theta_{k}-\eta_{\ell}\|_{2}$ for any $k\in[t]\backslash[K]$ and $\ell\in[s]\backslash[K]$, one can deduce that (2.8) holds. ∎
@@
$W_{1}\bigg{(}\sum_{k=1}^{t}p_{k}\delta_{z_{k}},\sum_{k=1}^{t}q_{k}\delta_{z_{k}}\bigg{)}\leq(z_{t}-z_{1})\cdot\max_{j\in[t-1]}\bigg{|}\sum_{k=1}^{j}(p_{k}-q_{k})\bigg{|}.$ (2.9)

### v6

@@ 4.2.1 Shape Correspondence
Figure 4(e) shows the target points $Y_{j}$’s colored by $s(j)$’s, where $s(j)\in\{1,\ldots,7\}$ is the index of the component that $Y_{j}$ belongs to. Now, we run Algorithm 1 and obtain the output $\pi\colon[n]\to[m]$. For each $X_{i}$, we verify if the part that it belongs to coincides with the part that its match $Y_{\pi(i)}$ belongs to, which is indexed by $s(\pi(i))$. To this end, Figure 4(b) shows the source points $X_{i}$’s colored by $s(\pi(i))$’s. We can observe that the proposed procedure recovers the corresponding regions mostly correctly. For instance, for $X_{i}$’s corresponding to the rostrum of the source dolphin, their colors $s(\pi(i))$’s (green) indicate that they are matched to the points corresponding to the rostrum of the target dolphin. Similarly, the points corresponding to the head (lime), tail (red), dorsal fin (blue), and peduncle (magenta) are mostly matched correctly, while there are some mismatches, for instance, among the points around the pectoral fin. Figure 4(c) and Figure 4(f) repeat the same procedure after removing 16 points around the fin of the target dolphin in Figure 4(e) and clustering again with k = $6$. Figure 4(c) shows that the matching result is similar to the previous one. In this case, though the points around the fin of the source dolphin do not have corresponding points in the target dolphin and thus we may ignore matching them, the proposed procedure still matches them to the points around the peduncle (magenta) of the target dolphin, which is a cluster adjacent to the removed fin.
Figure 5 repeats the same procedure for the 3D images of horses. Again, we can verify that the proposed procedure matches the corresponding body parts mostly correctly. Figure 5(c) and Figure 5(f) shows the results after removing 62 points around the two legs on the target horse’s left side in Figure 5(e) and clustering again with k = $6$. We again observe that the points in the common parts are mostly matched correctly.
(a) Source object
(b) Source points (color = $s(\pi(i))$)
(c) Source points (color = $s(\pi(i))$)
(d) Target object
(e) Target points (color = $s(j)$)
(f) Target points (color = $s(j)$)
Figure 5: 3D images of horses and the corresponding point clouds. (a) and (d) are the source and target 3D images, from which we obtain point clouds $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$, with $n=m=346$, as shown in (b) and (e), respectively. We apply k-means clustering to both point clouds with k =8=8 to cluster them based on the body parts. (e) shows the target points $Y_{j}$’s colored by $s(j)$’s, where $s(j)\in\{1,\ldots,8\}$ denotes the index of the component that $Y_{j}$ belongs to. (b) shows the source points $X_{i}$’s colored by $s(\pi(i))$’s, where $\pi$ is obtained by applying Algorithm 1. (c) and (f) repeat (b) and (e), respectively, after removing 62 points around the two legs on the target horse’s left side in (e) and clustering again with k =6=6.
4.2.2 Partial Matching for Alignment of Cryo-EM Structures
We apply the developed partial matching procedure to align cryogenic electron microscopy (cryo-EM) structures of human $\gamma$-secretase. Cryo-EM has gained significant attention in the past decade as a powerful tool for unveiling the structures of biomolecules at near-atomic resolution. However, many computational challenges arise; see [bendory2020single] for a review. Aligning cryo-EM structures is particularly challenging due to high noise levels and different conformations. Figure 6(a) shows the 3D model of human $\gamma$-secretase, which is thought to be involved in the development of Alzheimer’s disease. The cryo-EM data [^3] Listed as 5A63 at the Protein Data Bank (PDB) https://www.rcsb.org/structure/5A63 and as EMD-3061 at the Electron Microscopy Data Bank (EMDB) https://www.ebi.ac.uk/emdb/EMD-3061. of human $\gamma$-secretase [bai2015atomic] essentially represents a probability density function of the intensities of the object, stored as voxels on a 3D grid. Figures 6(b)-(d) show the projections of the cryo-EM data along the $x$, $y$, and $z$ axes, respectively, which clearly demonstrate high noise levels.
(a) 3D Model
(b) Projection ($x$-axis)
(c) Projection ($y$-axis)
(d) Projection ($z$-axis)
Figure 6: Human $\gamma$-secretase. (a) shows the 3D model of human $\gamma$-secretase, and (b)-(d) show the projections of the cryo-EM data along the $x$, $y$, and $z$ axes, respectively.
Alignment problems in cryo-EM are often tackled by using the density functions [singer2024alignment, harpaz2023three] with suitable transformations or by obtaining the point clouds from the density functions [riahi2023alignot, tajmir2025alignment], say, utilizing the topology representing network algorithm [martinetz1994topology]. The proposed algorithm can be applied to the latter setting. Here, we sample two point clouds from the cryo-EM data, one from the original density function and the other from the density function after applying a rigid transformation; see Figure 7(a). The goal is to recover the rigid transformation from the point clouds.
The proposed partial matching—Algorithm 1—provides correspondences between the points. Once we have the correspondences, we can estimate the rigid transformation using the standard orthogonal Procrustes problem—after centering—between matrices $\mathbb{X},\mathbb{Y}\in\mathbb{R}^{n\times d}$ whose rows are the point clouds $X_{1},\ldots,X_{n}$ and $Y_{\pi(1)},\ldots,Y_{\pi(n)}$, respectively, where $\pi$ is the output of Algorithm 1. Figure 7(b) shows the aligned point clouds, which seems reasonable. However, for better results, it is beneficial to subset the good matches by using the threshold $\rho$ in Algorithm 1. The crux is that we only need to partially match the points for alignment, and focusing on good matches can improve the alignment results. Figures 7(c) and 7(d) apply Algorithm 1 with $\rho$ being the $0.5$ and $0.1$ quantiles of $W(1,\pi(1)),\ldots,W(n,\pi(n))$ to use the best 50% and 10% of the correspondences, respectively. In these cases, we obtain the set of index $I$ from the output of Algorithm 1 and estimate the rigid transformation using the submatrices $\mathbb{X}_{I},\mathbb{Y}_{I}$ of $\mathbb{X},\mathbb{Y}$, respectively, corresponding to the rows indexed by $I$. Figures 7(c) and 7(d) indeed look better than using the full correspondences, which we can quantify by computing the Wasserstein-2 distance between the aligned point cloud and the target point cloud. The Wasserstein-2 distance between the aligned point cloud and the target point cloud is $10.68$ for the full correspondences, $9.04$ for the best 50%, and $8.31$ for the best 10%, which confirms that the partial matching improves the alignment results.
(a) Point clouds
(b) Aligned with full correspondences
(c) Aligned with best 50% correspondences
(d) Aligned with best 10% correspondences
Figure 7: Point clouds from the cryo-EM data of human $\gamma$-secretase and alignment results. (a) shows the point clouds sampled from the original density function and the density function after applying a rigid transformation, shown as source and target, respectively. For (b)-(c), we first apply Algorithm 1 to obtain the correspondences and then estimate the rigid transformation using the orthogonal Procrustes problem. (b) shows the aligned point clouds using all the correspondences, while (c) and (d) use the best 50% and 10% of the correspondences, respectively, based on the threshold $\rho$ in Algorithm 1.
We repeat the same process for the cryo-EM data of the voltage-gated calcium channel ($\text{Ca}_{\text{v}}1.1$) of rabbit [wu2016structure], which plays a role as the sensor for excitation-contraction coupling of skeletal muscles. Figure 8 shows the 3D model and projections of the cryo-EM data. Figure 9 shows the alignment results. Again, we obtain two point clouds from the cryo-EM data, one from the original density function and the other from the density function after applying a rigid transformation; see Figure 9(a). Figures 9(b)-(d) show the aligned point clouds using all the correspondences, the best 50%, and the best 10%, respectively. The Wasserstein-2 distance between the aligned point cloud and the target point cloud is $11.10$ for the full correspondences, $8.23$ for the best 50%, and $9.17$ for the best 10%. In this case, taking the best 50% correspondences yields the best partial matching result among the three settings.
(a) 3D Model
(b) Projection ($x$-axis)
(c) Projection ($y$-axis)
(d) Projection ($z$-axis)
Figure 8: Rabbit $\text{Ca}_{\text{v}}1.1$. (a) shows the 3D model of $\text{Ca}_{\text{v}}1.1$, and (b)-(d) show the projections of the cryo-EM data along the $x$, $y$, and $z$ axes, respectively.
(a) Point clouds
(b) Aligned with full correspondences
(c) Aligned with best 50% correspondences
(d) Aligned with best 10% correspondences
Figure 9: Point clouds from the cryo-EM data of rabbit $\text{Ca}_{\text{v}}1.1$ and alignment results. (a) shows the point clouds sampled from the original density function and the density function after applying a rigid transformation, shown as source and target, respectively. For (b)-(c), we first apply Algorithm 1 to obtain the correspondences and then estimate the rigid transformation using the orthogonal Procrustes problem. (b) shows the aligned point clouds using all the correspondences, while (c) and (d) use the best 50% and 10% of the correspondences, respectively, based on the threshold $\rho$ in Algorithm 1.
5 Conclusion
We have proposed matching procedures based on distance profiles, which are easily implementable without having to solve nontrivial optimization problems over permutations. Most importantly, we have shown that the proposed method is provably robust to outlier components and noise under some model assumptions that we believe capture essential aspects of empirical applications. To the best of our knowledge, this is the first work that provides theoretical guarantees for the robustness of matching procedures based on distance profiles in the context of object matching. Also, we have derived the new statistical complexity result of the distance profile distance—the lower bound to the Gromov-Wasserstein distance—that has provided insights into how to use the distance profile matching procedures in practice. We have also demonstrated the empirical performance of the proposed methods through simulations and applications to real data.
We suggest several directions for future research. On the theoretical side, it would be interesting to consider other types of matching guarantees in Theorem 3. As mentioned in Section 4, the matching accuracy (4.1) may be practically more useful. Similarly, for Theorem 4, it would be interesting to consider the probability that $\hat{\pi}$ is close to $\pi^{\ast}$ under a certain discrepancy, such as the Hamming distance, instead of the probability of $\hat{\pi}=\pi^{\ast}$. Also, one can explore a generalization based on mixtures beyond the Euclidean space, say, a general metric or Banach space. On the practical side, it would be interesting to delve deeper into the procedure of finding inliers, that is, the points from the common part of two objects. For interested readers, we discuss in Appendix C.2 how the threshold $\rho$ of Algorithm 1 can be chosen to identify the points from the common part at least in theory. However, developing a practical method applicable to real data is very challenging in general and needs further investigation. We leave these directions for future research.
Acknowledgments
Appendix A Proofs
A.1 Proof of Theorem 1
Proof.
where $\mathcal{F}_{\mathcal{X}}:=\bigcup_{x\in\mathcal{X}}\mathcal{F}_{x}=\{\phi_{x,r}:x\in\mathcal{X},r\geq 0\}$. By Theorem 8.23 of [vershynin_2018],
$\mathbb{E}\left[\sup_{x\in\mathcal{X}}W_{1}(\mu_{x},\hat{\mu}_{x})\right]\leq\mathbb{E}\left[\sup_{\phi\in\mathcal{F}_{\mathcal{X}}}|\mu\phi-\hat{\mu}\phi|\right]\leq C\sqrt{\frac{\mathrm{vc}(\mathcal{F}_{\mathcal{X}})}{n}}.$ (A.1)
where $C$ is an absolute constant that is independent of $x$, $n$, $\mathcal{X}$, and $\mu$. Here, $\mathrm{vc}(\mathcal{F}_{\mathcal{X}})$ is the Vapnik-Chervonenkis (VC) dimension of the function class $\mathcal{F}_{\mathcal{X}}$. As $\mathcal{X}\subset\mathbb{R}^{d}$, this is upper bounded by the VC dimension of the set of all closed Euclidean balls in $\mathbb{R}^{d}$, which is $d+1$; see [dudley1979balls]. Hence, we obtain (2.7). ∎
A.2 Proof of Proposition 1
Proof.
A.3 Proof of Proposition 2
Proof.
$W_{1}(\mu_{x},\mu_{x^{\prime}})=W_{1}\left(\sum_{k=1}^{t}p_{k}(\mu_{k})_{x},\sum_{k=1}^{t}p_{k}(\mu_{k})_{x^{\prime}}\right)\leq\sum_{k=1}^{t}p_{k}W_{1}((\mu_{k})_{x},(\mu_{k})_{x^{\prime}}),$ (A.2)
$\begin{split}W_{1}((\mu_{k})_{x},(\mu_{k})_{x^{\prime}})&=W_{1}\left((\mu_{k})_{\theta_{k}+\|x-\theta_{k}\|_{2}u},(\mu_{k})_{\theta_{k}+\|x^{\prime}-\theta_{k}\|_{2}u}\right)\\ &\leq\left\|\theta_{k}+\|x-\theta_{k}\|_{2}u-\theta_{k}-\|x^{\prime}-\theta_{k}\|_{2}u\right\|_{2}\\ &=\left|\|x-\theta_{k}\|_{2}-\|x^{\prime}-\theta_{k}\|_{2}\right|,\end{split}$ (A.3)
where the inequality uses Proposition 1. Combining (A.2) and (A.3), we obtain (2.8). ∎
A.4 Proof of Theorem 2
Proof.
[…]
Therefore,
[…]
where the first equality follows from the change of variables, and the second equality is due to (A.4). As $\Phi_{\#}\hat{\mu}$ is equivalent to the empirical measure of $\Phi_{\#}\mu$ which is defined on $\mathbb{R}^{t}$, we conclude from the standard results [dudley1969speed, fournier2015rate] that $\mathbb{E}(*)\leq W_{1}(\Phi_{\#}\mu,\Phi_{\#}\hat{\mu})\leq O(n^{-1/t})$. ∎
A.5 Proof of Theorem 3
@@
Meanwhile, notice that $X_{i}=\theta_{t(i)}+\xi_{i}$ and $Y_{j}=\eta_{s(j)}+\zeta_{j}$ for any $(i,j)\in[n]\times[m]$, where $\xi_{i}\,|\,t(i)$ and $\zeta_{j}\,|\,s(j)$ are independent sub-Gaussian random vectors. Recall that the Wasserstein-1 distance between two sets of points in $\mathbb{R}$ with the same cardinality satisfies the following:
@@
$\begin{split}&\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}W(i,j)-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}W(i,j)\\ &\geq\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}\overline{W}(t(i),s(j))-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}\overline{W}(t(i),s(j))-2R(Z_{1}+Z_{2})-2(U+V)\\ &\geq\min_{\beta\in[s]\backslash\{t(i)\}}\overline{W}(t(i),\beta)-\overline{W}(t(i),t(i))-2R(Z_{1}+Z_{2})-2(U+V)\\ &\geq\underbrace{\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)-R\cdot\bigg(1-\sum_{k\in[K]}p_{k}\bigg)}_{=:\Omega}-2R(Z_{1}+Z_{2})-2(U+V),\end{split}$ (A.5)
@@
Meanwhile, for $i\neq\ell$, note that $\xi_{i}-\xi_{\ell}\,|\,t(i),t(\ell)$ is a sub-Gaussian random vector with variance proxy $\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2}$. By Lemma 4, we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq\sum_{i\neq\ell}\mathbb{P}\left(\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}e^{2d}\exp\left(-\frac{(\Omega/8)^{2}}{16\sigma^{2}}\right).$
Similarly, we have
$\mathbb{P}\left(\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq m^{2}e^{2d}\exp\left(-\frac{(\Omega/8)^{2}}{16\tau^{2}}\right).$
@@
$\begin{split}\mathbb{P}(E)&\leq\mathbb{P}\left(Z_{1}\geq\frac{\Omega}{4R}\right)+\mathbb{P}\left(Z_{2}\geq\frac{\Omega}{4R}\right)+\mathbb{P}\left(U\geq\frac{\Omega}{8}\right)+\mathbb{P}\left(V\geq\frac{\Omega}{8}\right)\\ &\leq 2K(t+s-2)\exp\left(-\frac{2(n\wedge m)\Omega^{2}}{(8R)^{2}}\right)+2(n\vee m)^{2}e^{2d}\exp\left(-\frac{(\Omega/8)^{2}}{16\Gamma^{2}}\right)\\ &\leq\frac{\delta}{2}+\frac{\delta}{2}=\delta,\end{split}$
@@
[…]
To see this, it suffices to observe that we can replace $\Omega$ in the last line of (A.5) with $\Omega_{o}$, and the rest of the proof follows the same with $\Omega_{o}$ in place of $\Omega$. Since $\Omega_{o}\geq\Omega$, the condition (3.4) is a stronger condition than requiring $\Omega_{o}$ to be at least the right-hand side of (3.4). We chose to present (3.4) as it is written in terms of the outlier proportion which is more interpretable.
@@
$\overline{W}(\alpha,\alpha)\leq\bigg(1-\sum_{k\in[K]}p_{k}\bigg)\cdot\max_{(\alpha,\beta)\in[t]\times[s]}\|\theta_{\alpha}-\eta_{\beta}\|_{2}.$ (A.6)
@@
where $\lambda=\sum_{k\in[K]}p_{k}$. Since $|\|\theta_{\alpha}-\theta_{k}\|_{2}-\|\theta_{\alpha}-\eta_{\ell}\|_{2}|\leq\|\theta_{k}-\eta_{\ell}\|_{2}$ for any $k\in[t]\backslash[K]$ and $\ell\in[s]\backslash[K]$, one can deduce that (A.6) holds. ∎
@@
$W_{1}\bigg(\sum_{k=1}^{t}p_{k}\delta_{z_{k}},\sum_{k=1}^{t}q_{k}\delta_{z_{k}}\bigg)\leq(z_{t}-z_{1})\cdot\max_{j\in[t-1]}\bigg|\sum_{k=1}^{j}(p_{k}-q_{k})\bigg|.$ (A.7)